QUESTION 8

  1. According to the Coefficient Rule, in the expression: 5 x + 3 y <= 120, if y is increased by 1 unit while x is held constant, the quantity represented by 5 x + 3 y will increase by:

    3

    5

    x

    y

QUESTION 9

  1. Suppose each paperback box holds 4 paperback books and 3 hardcover books and each hardcover box holds 5 hardcover books and 2 paperback books. What are the coefficients in the constraint: __ P + __ H >= 15, representing total number of add l hardcovers >= 15?

    0, 5

    3, 0

    3, 5

    4, 2

QUESTION 10

  1. Suppose each paperback box holds 4 paperback books and 3 hardcover books and each hardcover box holds 5 hardcover books and 2 paperback books. What are the coefficients in the objective function: __ P + __ H, representing total number of add l books shipped?

    3, 5

    4, 2

    6, 8

    7, 7
Answer #1

Question 18.

According to the Coefficient Rule, in the expression: 5 x + 3 y< = 120, if y is increased by 1 unit while x is held constant, the quantity represented by 5 x + 3 y will increase by:

Correct answer is 3.

If we increase 1 unit of y to 5x + 3y then the revised expression is 5x + 3(y+1) = (5x + 3y) + 3. Hence, comparing earlier equation with the revised one, there is an increase of + 3.

Question 19.

Suppose each paperback box holds 4 paperback books and 3 hardcover books and each hardcover box holds 5 hardcover books and 2 paperback books. What are the coefficients in the constraint: __ P + __ H >= 15, representing total number of add l hardcovers> = 15?

For the expression __ P + __ H >= 15, total number is representing hardcovers. P represent paperback box and H represent hardcover box. Hence coefficients should represent number of hardcover books which can be accommodated in paperback box and hardcover box.

Hence, the expression is 3 P + 5 H >=15

The correct answer is 3,5.

QUESTION 10

Suppose each paperback box holds 4 paperback books and 3 hardcover books and each hardcover box holds 5 hardcover books and 2 paperback books. What are the coefficients in the objective function: __ P + __ H, representing total number of add l books shipped?

Since the objective function represent total number of books shipped. Hence, in this case P represent paperback books and H represent hardcover books.

Total paper back books in both boxes = 4 + 2 = 6

Total hardcover books in both boxes = 3 + 5 = 8

So, the objective function is 6 P + 8 H

The correct answer is 6,8.

If you liked my answer, please give like/ thumbs up. Thank you!